OpenStudy (anonymous):

Simplify the radical expression. sqrt of 7 (sqrt of 14 + sqrt of 3) please show the steps

OpenStudy (mathlegend):

\[\sqrt{7} + \sqrt{14} + \sqrt{3}\]

OpenStudy (mathlegend):

Is that what you have? @CalvinG

OpenStudy (anonymous):

Sorry just learning how to use the interface. It is \[\sqrt{7}\left( \sqrt{14}+ \sqrt{3} \right)\].

OpenStudy (mathlegend):

Oh, so I would distribute the radical 7... which means you must multiply what is inside of the radicals. So can you tell me what you get?

OpenStudy (anonymous):

So would it be \[\sqrt{7}\times \sqrt{14} and \sqrt{7}\times \sqrt{3}\]

OpenStudy (mathlegend):

Yes.

OpenStudy (mathlegend):

So what do you get?

OpenStudy (anonymous):

\[\sqrt{98}+\sqrt{21}\]

OpenStudy (mathlegend):

Correct!

OpenStudy (anonymous):

Is that the answer?

OpenStudy (mathlegend):

Now, we have to see if we can find a perfect square to go into 98 and then we can check if one goes into 21.

OpenStudy (anonymous):

ok

OpenStudy (mathlegend):

Nope, that is not the answer. I found a perfect square that goes into 98. So lets see if you can find it.

OpenStudy (anonymous):

\[7\times14\]

OpenStudy (mathlegend):

A perfect square as in ... 4 x 4 = 16 16 is the perfect square

OpenStudy (anonymous):

Ok I am lost. How would you go about finding the perfect square of 98.

OpenStudy (mathlegend):

Okay, so @CalvinG, what you want to do is... lets say for instance we used the perfect square 16. You would then check to see if 16 goes into 98. 98 / 16 = 6.125 (So we know that is not correct. Since the perfect square must go in it evenly... also remember that means we used 4 x 4) (So, try 5 x 5, 6 x 6, 7 x 7, etc... until one of the perfect squares goes into 98)

OpenStudy (anonymous):

\[7\times7=49\] which is halve of 98?

OpenStudy (mathlegend):

Correct!

OpenStudy (mathlegend):

So what we have now is..

OpenStudy (mathlegend):

\[\sqrt{49 * 2}\]

OpenStudy (mathlegend):

So take the perfect square and square root it. Once you get the square root then that number must come on the outside of the radical. Which would then leave the non perfect number to stay under the radical. So can you tell me what you get?

OpenStudy (anonymous):

\[7\sqrt{2}\]?

OpenStudy (mathlegend):

Yes!

OpenStudy (mathlegend):

Now we have.. \[7\sqrt{2} + \sqrt{21}\]

OpenStudy (mathlegend):

Can that 21 be broken down?

OpenStudy (anonymous):

\[7\times3\]

OpenStudy (mathlegend):

What number there is perfect?

OpenStudy (anonymous):

7

OpenStudy (mathlegend):

Nope.

OpenStudy (mathlegend):

3 x 3 = 9 9 is perfect

OpenStudy (anonymous):

So would it be \[3\sqrt{7}\]

OpenStudy (mathlegend):

Nope

OpenStudy (mathlegend):

Remember you are looking for something like... 2 x 2 = 4 4 is perfect... so you would need to check to see if the 4 goes into 21

OpenStudy (anonymous):

not perfectly

OpenStudy (mathlegend):

It must go into the 21 perfectly.

OpenStudy (anonymous):

Sorry evenly is the word i was looking for. lol

OpenStudy (anonymous):

so 21 can not be broken down?

OpenStudy (mathlegend):

Correct, 21 can not be broken. :)

OpenStudy (anonymous):

so is \[7\sqrt{2}+\sqrt{21}\] as far as it can be simplifyed?

OpenStudy (mathlegend):

Yup

OpenStudy (anonymous):

Thank you!